6
6 A charge of uniform volume density (40 ncum3) fills a cube with s.ocm is edges. What the total electric flux (in N.m2c) through the surface ofthis cube? a. 2.9 b. 2.0 c. 2.6 d. 2.3 e. 1.8
Answer #1

The Gaussian surface is the surface of the cube.

Volume of cube = 0.08m x0.08m x 0.08m = 5.12E-5m³

Total charge inside cube, q =5.12E-5 x40E-9 = 2.0E-11 C

Φ = q/ε₀ = 2.0E-11 /8.85E-12 = 2.3N m²/C (answer d))